Could you please help me out with this problem??
Factor completely...
7w^3+w^6-30
Thanks for your help!

Answer by lazi(15)

--30
-/--\
3 --10
10 - 3 = 7
w^6 + 7w^3 -30 = 0
(w^3 + 10)(w^3 - 3) = 0